# Aperiodic Structures Never Collapse: Fibonacci Hierarchies for Lossless Compression

### 16 Mar 2026 v1

@@ 5.3 Implementation
Quasicryth v5.5 is implemented in ANSI C99. The implementation uses 24-bit arithmetic coding with range coding, bzip2 via libbz2, and zlib for codebook compression. No SIMD or hardware-specific optimisations are used. The codebase is approximately 4,000 lines of C. All experiments were run on a single core; no multi-threading is employed.
@@ 6.1 Compression Performance
Table 5 reports compression ratios and sizes. On enwik9, Quasicryth achieves 35.99% (359,883,431 B), with the full breakdown in Table 6.
Table 5: Compression ratios (compressed/original $\times 100\%$) for Quasicryth v5.5 versus standard compressors.
@@
Quasicryth v5.5 36.92 39.29 38.51 37.70 35.99
gzip -9 35.63 36.29 36.85 36.45 32.26
bzip2 -9 28.41 28.89 29.16 29.01 25.40
xz -9 31.89 28.39 27.21 24.87 21.57
@@
Payload (AC) 254,952,735
Escapes (bz2) 83,537,358
Codebook (zlib) 77,585
Case flags (AC) 21,315,698
Total 359,883,431
Ratio 35.99%
@@ 6.2 Deep Hierarchy Hit Counts
Table 7: Deep hierarchy hits by level and file size. All counts are from the Fibonacci tiling; Period-5 achieves zero at every level shown (hierarchy collapsed). Levels 8–9 (89g, 144g) first activate at enwik9 scale.
File Words 13g 21g 34g 55g 89g 144g Total deep
alice29.txt 36K 2 0 0 0 0 0 2
enwik8_3M 821K 971 195 77 0 0 0 1,243
enwik8_10M 2.8M 2,715 383 140 47 0 0 3,285
enwik8 27.7M 24,910 3,256 1,369 551 0 0 30,086
enwik9 298.3M 652,124 109,492 40,475 7,222 2,396 945 812,654
@@
Figure 4: Deep hierarchy hits (levels 4–9, Fibonacci-exclusive) per word versus corpus size (log-log). Labels show absolute hit counts. The ratio is roughly flat from enwik8_3M to enwik8 while fixed-size codebooks limit density; at enwik9, levels 8–9 activate and the ratio jumps to $2.7\times 10^{-3}$, confirming new $O(W)$ terms.
152 KB3 MB10 MB100 MB1 GB0%5050%100100%Word-weighted share of deep hits13-gram21-gram34-gram55-gram89-gram144-gram
@@ 6.3 Timing
Table 8: Compression and decompression times (C implementation, single core).
File Size Compress Decomp. C/D
alice29.txt 152 KB 0.1s <<0.1s
[…]
@@ 6.4 Asymmetric Compression Profile
Why compression is expensive. Compression performs three costly operations: (1) the phase search evaluates 32 candidate tilings with the full substitution hierarchy, scoring each against the codebooks; (2) codebook construction counts all n-gram frequencies up to 144 words per entry, with periodic pruning passes for the 89-gram and 144-gram levels at the 10 M+ word scale; (3) the 32-phase scoring loop must build the complete hierarchy at each candidate phase, repeating the $O(W)$ hierarchy detection.
Why decompression is cheap. The decompressor reads the 2-byte phase from the header and regenerates the entire tiling deterministically — no search, no n-gram counting, no frequency tables. It then decompresses the zlib’d codebook (77 KB at enwik9 scale), decompresses the bz2 escape buffer, and streams through the AC payload once, consulting preloaded codebook hash tables. The entire decompression is a single sequential pass.
Growing asymmetry at scale. The C/D ratio increases from $6.3\times$ at 3 MB to $15.9\times$ at 1 GB (Figure 6), because the 89-gram and 144-gram frequency counting — which hashes windows of 89–144 consecutive word IDs — dominates codebook construction at the enwik9 scale. Decompression throughput remains roughly constant at $\sim$10 MB/s regardless of file size, while compression throughput falls to $\sim$0.65 MB/s at 1 GB.
Practical implication. For any file decompressed $k$ times, the amortised cost per access is $T_{C}/k+T_{D}$. At enwik9 scale ($T_{C}=1{,}540$ s, $T_{D}=97$ s), even a single decompression amortises the compression cost within $\sim$16 accesses. The upfront compression investment is a one-time payment; every subsequent decompression is fast.
152 KB3 MB10 MB100 MB1 GB055101015152020556.36.311.511.514.214.215.915.9Compress / Decompress ratio
Figure 6: Compression-to-decompression time ratio (C/D ratio) versus file size. The ratio grows from $6.3\times$ at 3 MB to $15.9\times$ at 1 GB as the 89-gram and 144-gram frequency counting begins to dominate codebook construction. Decompression throughput remains stable at $\sim$10 MB/s across all scales; the growing asymmetry is entirely a compression-side effect.
@@ 7.1 QC Contribution vs All-Unigram Baseline
Table 9 reports the payload savings of the Fibonacci tiling over an all-unigram baseline (every tile is $S$, no bigram or n-gram attempts), using identical codebooks and escape streams. This isolates the compression value of the entire multi-level parsing structure.
Table 9: Quasicrystal contribution: payload savings over all-unigram baseline.
@@
alice29.txt 152 KB 36.92% 1,276 B 0.84%
enwik8_3M 3 MB 39.29% 94,321 B 3.14%
enwik8_10M 10 MB 38.51% 312,638 B 3.13%
enwik8 100 MB 37.70% 2,838,937 B 2.84%
enwik9 1,000 MB 35.99% 45,608,715 B 4.56%
The QC contribution increases at larger scales, reaching 4.56% of the original file size at enwik9. This growing advantage reflects the activation of deep hierarchy levels with increasing input size.
@@ 7.2 Aperiodic Advantage: Fibonacci vs Period-5
enwik8_3M 895,581 802,632 801,260
enwik8_10M 2,966,773 2,659,942 2,654,135
enwik8 29,036,210 26,237,097 26,196,712
enwik9 300,561,450 256,302,106 254,952,735
@@
enwik8_3M 1,372 B 92,949 B
enwik8_10M 5,807 B 306,831 B
enwik8 40,385 B 2,799,113 B
enwik9 1,349,371 B 44,259,344 B
@@
152 KB3 MB10 MB100 MB1 GB0%22%44%66%File size (MB)QC savings (% of original)
Figure 7: QC tiling contribution (payload savings over all-unigram baseline, as % of original) versus file size, semi-log scale. The contribution rises to 4.56% at 1 GB as deeper hierarchy levels progressively activate.
@@ 7.4 Approaches Rejected
  • Word-level LZ pre-pass. Finding and encoding repeated word sequences as back-references improved overall compression but destroyed the QC contribution (reduced to 334 B on alice29) because the LZ layer captures the same long-range redundancy exploited by the bigram/n-gram codebooks; the two mechanisms compete.
@@ 8. Conclusion
We have presented Quasicryth v5.5, a lossless text compressor based on a deep Fibonacci quasicrystal substitution hierarchy, and proved that this hierarchy never collapses while all periodic alternatives do. The proof synthesises Perron-Frobenius analysis of the substitution matrix, the Pisot-Vijayaraghavan property of $\varphi$, Sturmian balance and aperiodicity, and Weyl’s equidistribution theorem. These are not independent observations: they are aspects of a single mathematical object — the Fibonacci quasicrystal — that jointly guarantee the compression properties of Quasicryth.
The aperiodic advantage is proven and measurable: 1,349,371 B at enwik9 (1 GB), growing superlinearly because new hierarchy levels (89-gram, 144-gram) activate at scale. At enwik9, 812,654 deep hierarchy positions (levels 4–9) are exploited; Period-5, the best periodic approximant with the same L/S ratio, achieves zero deep positions, because its hierarchy collapses at level 4 — exactly as the theory predicts.

### 23 Mar 2026 v2

@@ 5.3 Implementation
Quasicryth v5.6 is implemented in ANSI C99. The implementation uses 24-bit arithmetic coding with range coding, LZMA via liblzma for escape and codebook compression. No SIMD or hardware-specific optimisations are used. The codebase is approximately 4,000 lines of C. All experiments were run on a single core; no multi-threading is employed.
@@ 6.1 Compression Performance
Table 5 reports compression ratios and sizes. On enwik8, Quasicryth achieves 26,247,496 B (26.25%) in multi-structure mode and 27,026,429 B (27.03%) in Fibonacci-only mode. On enwik9, Quasicryth achieves 225,918,349 B (22.59%) in multi-structure mode and 234,560,637 B (23.46%) in Fibonacci-only mode. Quasicryth now surpasses bzip2 on all benchmarks and approaches xz. The full breakdown is in Table 6.
Table 5: Compression ratios (compressed/original $\times 100\%$) for Quasicryth v5.6 versus standard compressors.
@@
Quasicryth v5.6 (multi) 35.60 31.85 29.83 26.25 22.59
Quasicryth v5.6 (fib) 35.91 32.55 30.56 27.03 23.46
gzip -9 35.63 36.28 36.85 36.44 32.26
bzip2 -9 28.40 28.88 29.16 29.00 25.40
xz -9 31.88 28.38 27.21 24.86 21.57
@@
Payload (AC) 180,760,315
Escapes (LZMA) 24,227,220
Codebook (LZMA) 533,680
Case flags (AC) 20,397,073
Total 225,918,349
Ratio 22.59%
@@ 6.2 Deep Hierarchy Hit Counts
Table 7: Deep hierarchy hits by level and file size: Fibonacci-only (F) vs Multi-structure (M). Period-5 achieves zero at every level shown (hierarchy collapsed). Levels 8–9 (89g, 144g) first activate at enwik9 scale. Multi-structure tilings significantly increase deep hits: 55-gram ++56%, 34-gram ++20%, 21-gram ++18% on enwik8, driven by the optimized alpha set including $\alpha=0.502$ (far below golden ratio) which contributes massive trigram/5-gram coverage. Since deeper hits cover exponentially more words per symbol, this redistribution drives the payload reduction (Table 15).
13g 21g 34g 55g 89g 144g Total deep
File Words F M F M F M F M F M F M F M
alice29.txt 36K 9 9 0 0 0 0 0 0 0 0 0 0 9 9
enwik8_3M 821K 2,454 2,469 405 489 98 98 39 65 0 0 0 0 2,996 3,121
enwik8_10M 2.8M 9,460 9,400 1,245 1,553 295 346 83 131 8 8 5 5 11,096 11,443
enwik8 27.7M 89,615 87,344 13,344 15,736 3,434 4,118 936 1,464 216 224 85 85 107,630 108,971
enwik9 298.3M 1,998,119 1,890,784 372,974 424,084 112,599 153,713 25,277 36,776 5,369 5,544 2,026 2,026 2,516,364 2,512,927
@@
Per-Family Tiling Coverage
Table 7 reveals that multi-structure tilings redistribute deep hits toward deeper levels: 55-gram counts increase by 56%, 34-gram by 20%, and 21-gram by 18% on enwik8, while 13-gram counts decrease slightly, as the additional tilings (especially optimized $\alpha=0.502$) find deeper matches at positions where Fibonacci-only found only 13-grams. Since a 55-gram hit encodes 55 words with one AC symbol versus 4.2 average words for a 13-gram, this redistribution substantially improves coding efficiency. The multi-structure advantage also manifests in tiling coverage: the number of word positions at which at least one tiling produces a deep hierarchy entry point. Table 8 decomposes this coverage by tiling family.
Table 8: Per-family tiling contribution to deep positions. Each row shows cumulative deep positions as tiling families are added. The golden-ratio (Fibonacci) family provides the base; 6 original non-golden tilings add 19%, and 18 optimized additions (including $\alpha=0.502$) add a further 2.6%, for 36 tilings total. Both enwik8 and enwik9 show the same +22.0% gain.
Tiling family enwik8 deep pos. Cumul. enwik9 deep pos. Cumul.
Golden ($1/\varphi$, 12 tilings) 9,777,069 9,777,069 132,072,408 132,072,408
++\;Orig. non-golden (6) ++1,854,993 11,632,062 ++25,053,269 157,125,677
++\;Optimized (18) ++296,848 11,928,910 ++3,996,271 161,121,948
Total (36 tilings) 11,928,910 161,121,948
Gain over Golden-only +22.0% +22.0%
Table 8 shows that the 12 golden-ratio tilings provide the majority of deep positions (9,777,069 on enwik8, 132,072,408 on enwik9). The 6 original non-golden tilings add 19.0% (1,854,993 positions on enwik8), while the 18 optimized additions discovered by greedy alpha search contribute a further 2.6% (296,848 positions on enwik8).
Table 9 provides a detailed per-tiling breakdown of new codebook-matched positions at each n-gram level on enwik9. The golden-ratio family is the only one reaching levels 8–9 (89-gram, 144-gram). The $\alpha=0.502$ tiling contributes 2.7M new trigram positions but zero hits at 13-gram or above — its value is pure shallow-level volume at positions unreachable by near-golden tilings. In contrast, $\alpha=0.619$ is the most “deep-capable” optimized alpha, finding matches up to 89-gram level, with strong contributions across 5-gram through 34-gram.
Table 9: Per-tiling new codebook-matched positions by n-gram level (enwik9, 298.3M words). Each row shows the marginal contribution of that tiling pair after all preceding tilings have been applied. The golden-ratio family exclusively provides 89-gram and 144-gram coverage; $\alpha=0.502$ contributes only at the trigram level; near-golden optimized alphas fill 5-gram through 34-gram gaps.
# Tiling ($\alpha$) Tilings tri 5g 8g 13g 21g 34g 55g 89g 144g Total new
0–11 Golden ($1/\varphi=0.618$) 12 132,071K 62,073K 27,322K 13,147K 3,048K 1,009K 233K 68K 20K 238,991K
12–13 $\sqrt{58}{-}7$ (0.616) 2 13,827K 6,606K 2,503K 840K 396K 18K 8K 24,198K
14–15 noble-5 (0.612) 2 7,359K 4,597K 2,170K 598K 304K 15,028K
16–17 $\sqrt{13}{-}3$ (0.606) 2 3,867K 3,109K 1,954K 248K 126K 9,303K
18–19 opt-0.502 2 2,655K 46K 35K 2,736K
20–21 opt-0.619 2 635K 2,548K 1,161K 811K 208K 170K 1.5K 0.8K 5,535K
22–23 opt-0.617 2 338K 1,771K 997K 640K 261K 71K 32K 4,110K
24–25 opt-0.616 2 179K 1,249K 829K 539K 270K 26K 12K 3,105K
26–27 opt-0.620 2 93K 920K 619K 606K 136K 131K 2,504K
28–29 opt-0.614 2 50K 614K 579K 373K 229K 1,845K
30–31 opt-0.621 2 25K 466K 403K 509K 95K 99K 1,598K
32–33 opt-0.622 2 13K 328K 326K 464K 68K 72K 1,272K
34–35 opt-0.612 2 8K 214K 336K 220K 167K 943K
Total (36 tilings) 36 161,121K 84,540K 39,235K 18,993K 5,307K 1,598K 287K 69K 20K 311,170K
Figure 4 shows how the greedy-selected level distribution changes as tiling families are added. Each line shows a level’s greedy-selected count as a percentage of its golden-ratio baseline: values above 100% indicate growth, while decreasing slopes indicate redistribution to deeper levels. The trigram line peaks at $\alpha=0.502$ (119%) then declines to 109% as near-golden alphas upgrade those positions; 21-gram grows to 181% and 34-gram to 169% — confirming that the optimized alphas redistribute shallow matches to deeper, more efficient encodings.
Golden+sq58+nob5+sq13+o502+o619+o617+o616+o620+o614+o621+o622+o612100100%120120%140140%160160%180180%100100%Tiling group (cumulative)% of golden-ratio baselinetri5g8g13g21g34g55g
Figure 4: Greedy-selected level distribution as tiling families are added (enwik9), normalised to golden-ratio baseline (= 100%). Deeper levels (21g, 34g) grow most aggressively — up to 181% and 169% of their golden baseline — as optimized alphas upgrade positions from trigram/5-gram to deeper matches. The trigram line peaks at $\alpha=0.502$ (119%, pure tri volume) then declines as subsequent near-golden alphas redistribute those positions upward. 89-gram and 144-gram (not shown) remain at $\approx$100% (golden-exclusive).
@@
Figure 5: Deep hierarchy hits (levels 4–9, Fibonacci-exclusive) per word versus corpus size (log-log). Labels show absolute hit counts. The ratio is roughly flat from enwik8_3M to enwik8 ($\approx 4\times 10^{-3}$); at enwik9, levels 8–9 activate and the ratio jumps to $8.4\times 10^{-3}$, confirming new $O(W)$ terms.
152 KB3 MB10 MB100 MB1 GB0%5050%100100%Word-weighted shareFibonacci-only (12 tilings)13g21g34g55g89g144g
152 KB3 MB10 MB100 MB1 GB0%5050%100100%Word-weighted shareMulti-structure (36 tilings)13g21g34g55g89g144g
@@
152 KB3 MB10 MB100 MB1 GB2020%2525%3030%3535%4040%File size (MB)Compression ratio (%)Quasicryth Multi (36)Quasicryth Fib (12)gzip -9bzip2 -9xz -9
Figure 7: Compression ratio versus file size (semi-log scale). Quasicryth improves steadily with corpus size, crossing bzip2 at $\approx$500 KB and narrowing the gap to xz from 3.7 $p$ at 152 KB to 1.0 $p$ at 1 GB. Multi-structure tilings (solid) consistently outperform Fibonacci-only (dashed) by 0.31–0.87 pp. gzip’s byte-level LZ77 is relatively size-insensitive; Quasicryth’s word-level codebooks and deep hierarchy benefit disproportionately from larger corpora.
@@ 6.3 Timing
Table 10: Compression and decompression times (C implementation, single core). Format: Fib-only / Multi-struct.
File Size Compress (Fib / Multi) Decomp. (Fib / Multi)
alice29.txt 152 KB 0.09s / 0.10s 0.01s / 0.01s
enwik8_3M 3 MB 1.46s / 2.05s 0.14s / 0.14s
enwik8_10M 10 MB 8.41s / 11.80s 0.48s / 0.47s
enwik8 100 MB 82.88s / 120.32s 4.84s / 4.74s
enwik9 1,000 MB 1,070s / 1,476s 46s / 45s
@@ 6.4 Asymmetric Compression Profile
Why compression is expensive. Compression performs three costly operations: (1) the phase search evaluates 36 candidate tilings with the full substitution hierarchy, scoring each against the codebooks; (2) codebook construction counts all n-gram frequencies up to 144 words per entry, with periodic pruning passes for the 89-gram and 144-gram levels at the 10 M+ word scale; (3) the 36-tiling scoring loop must build the complete hierarchy at each candidate phase, repeating the $O(W)$ hierarchy detection.
Why decompression is cheap. The decompressor reads the 2-byte phase from the header and regenerates the entire tiling deterministically — no search, no n-gram counting, no frequency tables. It then decompresses the LZMA’d codebook (534 KB at enwik9 scale), decompresses the LZMA escape buffer, and streams through the AC payload once, consulting preloaded codebook hash tables. The entire decompression is a single sequential pass.
Growing asymmetry at scale. The C/D ratio increases from $15\times$ at 3 MB to $33\times$ at 1 GB (Figure 8), because the 89-gram and 144-gram frequency counting — which hashes windows of 89–144 consecutive word IDs — dominates codebook construction at the enwik9 scale. Decompression throughput remains roughly constant at $\sim$22 MB/s regardless of file size, while compression throughput falls to $\sim$0.68 MB/s at 1 GB.
Practical implication. For any file decompressed $k$ times, the amortised cost per access is $T_{C}/k+T_{D}$. At enwik9 scale ($T_{C}=1{,}476$ s, $T_{D}=45$ s), even a single decompression amortises the compression cost within $\sim$33 accesses. The upfront compression investment is a one-time payment; every subsequent decompression is fast.
152 KB3 MB10 MB100 MB1 GB0101020203030101014.614.625.125.125.425.432.932.9Compress / Decompress ratio
Figure 8: Compression-to-decompression time ratio (C/D ratio) versus file size. The ratio grows from $15\times$ at 3 MB to $33\times$ at 1 GB as the 89-gram and 144-gram frequency counting begins to dominate codebook construction. Decompression throughput remains stable at $\sim$22 MB/s across all scales; the growing asymmetry is entirely a compression-side effect.
@@ 7.1 QC Contribution vs All-Unigram Baseline
Table 11 reports the total compressed size savings of the Fibonacci tiling over a no-tiling baseline, using identical codebooks and escape streams. This isolates the compression value of the entire multi-level parsing structure.
Table 11: Quasicrystal contribution: Fibonacci vs No-tiling (savings in total compressed size).
@@
alice29.txt 152 KB 35.91% 1,971 B 1.30%
enwik8_3M 3 MB 32.55% 37,098 B 1.24%
enwik8_10M 10 MB 30.56% 148,350 B 1.48%
enwik8 100 MB 27.03% 1,756,822 B 1.76%
enwik9 1,000 MB 23.46% 20,735,733 B 2.07%
The QC contribution increases at larger scales, reaching 2.07% of the original file size at enwik9. This growing advantage reflects the activation of deep hierarchy levels with increasing input size.
@@ 7.2 Aperiodic Advantage: Fibonacci vs Period-5
enwik8_3M 685,441 684,586 648,343
enwik8_10M 2,373,919 2,351,600 2,225,569
enwik8 23,544,808 23,039,959 21,787,986
enwik9 210,138,336 200,492,072 189,402,603
@@
enwik8_3M 36,243 B 855 B
enwik8_10M 126,031 B 22,319 B
enwik8 1,251,973 B 504,849 B
enwik9 11,089,469 B 9,646,264 B
Table 14: Multi-structure advantage: Multi-tiling over Fibonacci-only (bytes saved in total compressed size).
File Multi saves over Fib
alice29.txt 458 B
enwik8_3M 20,723 B
enwik8_10M 73,276 B
enwik8 778,933 B
enwik9 8,642,288 B
Table 15: Compressed file breakdown: Fibonacci-only vs Multi-structure.
enwik8 (100 MB) enwik9 (1 GB)
Component Fib Multi Fib Multi
Payload (AC) 21,787,986 21,009,053 189,402,603 180,760,315
Escapes (LZMA) 2,800,956 2,800,956 24,227,220 24,227,220
Codebook (LZMA) 541,664 541,664 533,680 533,680
Case (AC) 1,895,762 1,895,762 20,397,073 20,397,073
Total 27,026,429 26,247,496 234,560,637 225,918,349
Ratio 27.03% 26.25% 23.46% 22.59%
Multi saves 778,933 B (-0.78 pp) 8,642,288 B (-0.87 $p$)
Table 15 reveals that the multi-structure advantage is concentrated entirely in the payload stream: the escape stream, codebook, and case flags are identical between Fibonacci-only and multi-structure modes, because they depend on the codebook and input text, not on the tiling. Only the arithmetic-coded payload differs, where the additional non-Fibonacci tilings provide complementary deep hierarchy positions that improve n-gram hit rates. At enwik8 scale, the payload shrinks by 778,933 B (-3.57%); at enwik9, by 8,642,288 B (-4.56%).
@@
152 KB3 MB10 MB100 MB1 GB0%11%22%33%File size (MB)QC savings (% of original)
Figure 9: QC tiling contribution (Fibonacci savings over no-tiling, as % of original) versus file size, semi-log scale. The contribution rises to 2.07% at 1 GB as deeper hierarchy levels progressively activate.
@@ 7.4 Approaches Rejected
  • Byte-split indices. Splitting codebook indices into high/low byte streams for separate entropy coding added overhead without improving compression; the AC model already captures inter-symbol dependencies efficiently.
  • Semi-static two-pass. A first pass to gather statistics followed by a second encoding pass doubled compression time but yielded negligible improvement over the single-pass adaptive model.
  • Full-vocab unigrams. Encoding the entire vocabulary as unigrams (no bigram or n-gram codebooks) reduced codebook overhead but destroyed the hierarchical parsing advantage entirely.
  • Multi-stream LZMA payload. Splitting the AC payload into multiple LZMA-compressed streams by context type added framing overhead that exceeded the marginal compression gain from specialised contexts.
  • Model mixing. PAQ-style adaptive mixing of multiple context models (unigram, bigram, hierarchy-based) increased compression time by $5\times$ while improving the ratio by less than 0.1%; the single adaptive model proved sufficient.
@@ 8. Conclusion
We have presented Quasicryth v5.6, a lossless text compressor based on a deep Fibonacci quasicrystal substitution hierarchy with 36 multi-structure tilings, and proved that this hierarchy never collapses while all periodic alternatives do. The proof synthesises Perron-Frobenius analysis of the substitution matrix, the Pisot-Vijayaraghavan property of $\varphi$, Sturmian balance and aperiodicity, and Weyl’s equidistribution theorem. These are not independent observations: they are aspects of a single mathematical object — the Fibonacci quasicrystal — that jointly guarantee the compression properties of Quasicryth.
The aperiodic advantage is proven and measurable: 11,089,469 B at enwik9 (1 GB), growing superlinearly because new hierarchy levels (89-gram, 144-gram) activate at scale. At enwik9, 2,512,927 deep hierarchy positions (levels 4–9) are exploited; Period-5, the best periodic approximant with the same L/S ratio, achieves zero deep positions, because its hierarchy collapses at level 4 — exactly as the theory predicts.
The optimized alpha selection process — an iterative greedy search over candidate irrational slopes — discovered that $\alpha=0.502$, far below the golden ratio $1/\varphi\approx 0.618$, provides massive complementary trigram and 5-gram coverage. The resulting 36-tiling engine (12 golden-ratio, 6 original non-golden, 18 optimized additions) achieves 26.25% on enwik8, a 0.78 pp improvement over Fibonacci-only.
With multi-structure tilings, LZ77 word-level pre-pass, LZMA escape compression, and an improved variable-context model, Quasicryth v5.6 now surpasses bzip2 on all benchmarks and approaches xz compression ratios.